which points are on the graph of f(x) = 10^x
Choose al answers that are correct.
(-2,1/100)
(-1,1/10)
(0,10)
(2,20)

Answer by ewatrrr(24785)

f(x) = 10^x
Choose al answers that are correct.
(-2,1/100)
(-1,1/10)
..........Note
10^0 = 1
10^2 = 200